A screw gauge has 50 divisions on its circular scale.The pitch is 0.5mm.Find its least count?

Grade:8

1 Answers

Arun
5 years ago
least count=pitch / no. of divisions on circular scale=0.5/50=1/100=0.01 mm
 
 
Regards
Arun (askIITians forum expert)